# CodeHalu: Investigating Code Hallucinations in LLMs via Execution-based Verification

### v3

@@ Abstract
CodeHalu: Investigating Code Hallucinations in LLMs via
Execution-based Verification
Yuchen Tian1thanks:   Equal contribution. Corresponding to: weixiangyan@ucsb.edu     Weixiang Yan211footnotemark: 1     Qian Yang3,4     Xuandong Zhao5     Qian Chen6 Wen Wang6     Ziyang Luo7     Lei Ma1     Dawn Song5 1The University of Tokyo  2UC, Santa Barbara  3Mila-Québec AI Institute  4Université de Montréal 5UC, Berkeley   6Alibaba Group   7Hong Kong Baptist University yuchentovo@gmail.com      weixiangyan@ucsb.edu      qian.yang@mila.quebec xuandongzhao@berkeley.edu      {tanqing.cq,w.wang}@alibaba-inc.com cszyluo@comp.hkbu.edu.hk      ma.lei@acm.org      dawnsong@cs.berkeley.edu
1 Introduction
@@
Deep neural networks often generate erroneous information that contradicts the original content, cannot be verified, or conflicts with real-world knowledge. This phenomenon, commonly known as model hallucination, attracts widespread attention in the fields of natural language processing and multimodal learning (Ji et al., 2023; Zhang et al., 2023; Liu et al., 2024), with the community actively exploring methods to mitigate hallucinations (Peng et al., 2023; Elaraby et al., 2023; Liu et al., 2023). However, the issue of model hallucination in the code generation domain remains unexplored.
Conducting a thorough and dedicated study on code hallucinations is crucial for improving the quality of code generated by LLMs. Firstly, the purpose of code is to solve problems, and its value is realized only when the code executes successfully and passes tests (Chen et al., 2021; Austin et al., 2021; Yan et al., 2023). This necessitates that the generated code not only maintain strict logic and precision but also undergoes execution verification to confirm its correctness. Therefore, the practical use and verification of code differ significantly from Natural Language(NL) texts, meaning we cannot directly apply the definitions and methods used for NL hallucinations to code. Secondly, code snippets containing hallucinations may trigger runtime errors, or exhibit functional defects, which hinder the reliable deployment of LLMs in automated software development scenarios. Lastly, by exploring and verifying code hallucinations in a targeted manner, we can effectively uncover their causes and contribute to improving the architecture and training methods of LLMs.
To fill this gap, we define the concept of code hallucination in LLMs, based on the unique purpose and function of the code. Code hallucinations refer to the phenomenon where code generated by LLMs is syntactically correct or even semantically plausible but ultimately cannot execute as expected or fails to meet specified requirements. [^1] We test 16 LLMs using 105,958 code samples. The experimental results demonstrate that only 9 models occasionally exhibit syntactic errors in the generated code, with an exceptionally low average error rate of 0.0020. These findings support our initial hypothesis that the code generated by LLMs is generally syntactically correct and even semantically plausible or appropriate. Detailed statistical data and experimental results are shown in Appendix A. This phenomenon typically arises from various factors, such as errors or outdated information in the training data, an inadequate grasp of the syntax rules and programming paradigms of the programming languages, and limitations in the logical processing capabilities of the models. In contrast to previous methods that passively explore hallucinations in NLP through a Q&A framework or by prompting LLMs to generate hallucinated answers (Lin et al., 2021; Cheng et al., 2023), we employ an active strategy to detect hallucinations during the code generation process by LLMs. This approach is crucial as the ultimate goal of the generated code is to execute correctly and fulfill specific tasks.
@@ 2 Related Work
In the field of NLP, hallucination is initially defined as the phenomenon where the text generated by a model is fluent and natural but either lacks substantive meaning or is inconsistent with the provided source content (Ji et al., 2023). Recently, Zhang et al. (2023) standardize the definition of NL hallucinations in LLMs into three categories: input-conflicting hallucinations, where the content generated by LLMs diverges from the user’s input; context-conflicting hallucinations, in which the generated content contradicts previously generated content; and fact-conflicting hallucinations, where the generated content conflicts with established world knowledge. These hallucinations are attributed to various factors, such as poor-quality data samples in the training dataset or the use of sampling algorithms with high uncertainty.
In the multimodal domain, Zhai et al. (2023) classify types of hallucinations in image-to-text scenarios, such as image captioning and visual question answering. They define three main types of hallucinations: object existence hallucinations, object attribute hallucinations, and object relationship hallucinations. In text-to-image scenarios, such as image generation, hallucinations refer to the creation of factually incorrect details by the image generation model in response to the given text input. Huang et al. (2024) introduce VHTest, which evaluates hallucinations across eight dimensions in images, including the existence, shape, color, orientation, OCR, size, position, and counting of visual objects. In text-to-video scenarios, such as video generation, Chu et al. (2024) define three types of hallucinations: prompt consistency hallucinations, static hallucinations, and dynamic hallucinations. Detailed definitions and descriptions of each type of hallucination are provided in Appendix A. Although the issue of hallucinations receives extensive attention in NLP and multimodal domains, it remains unexplored in the code domain. Therefore, we propose CodeHalu to systematically define, identify, classify, and quantify code hallucinations in LLMs.
@@
In recent years, numerous studies focus on evaluating the capability of LLMs to handle various programming tasks. Among these, the HumanEval (Chen et al., 2021), includes 164 Python programming problems, each with an average of 6.7 unit tests. The MBPP (Austin et al., 2021) benchmark contains 974 Python programming tasks. Compared to these, the APPS (Hendrycks et al., 2021) benchmark presents more challenging programming questions, with each problem averaging 293.2 words in length. CodeScope (Yan et al., 2023) covers 43 programming languages and eight coding tasks to comprehensively evaluate LLMs in code understanding and generation. MMCode (Li et al., 2024) is designed to evaluate the programming capability of code models in multimodal scenarios. SWE-bench (Jimenez et al., 2023) evaluates the capability of LLMs to edit code repositories to solve problems with a level of complexity similar to that faced by human programmers in real-world programming situations. Overall, existing code benchmarks focus on evaluating the performance of LLMs on various programming tasks. However, there is still a lack of effective methods to detect and quantify potential hallucinations that may occur in code generation. Therefore, we propose CodeHaluEval to detect and quantify code hallucinations in LLMs.
3 Code Hallucination
@@
In multiple domains, existing work (Ji et al., 2023; Zhang et al., 2023; Huang et al., 2024; Zhai et al., 2023; Chu et al., 2024) often equates errors with hallucinations, or considers errors as a specific subset of hallucinations. We follow this perspective and regard code errors as a specific subset of code hallucinations. In other words, errors manifest as a form of hallucination, but not all hallucinations can be adequately expressed through errors. Figure 2 illustrates the distinction between code errors and code hallucinations. The code on the left exhibits a typical code error due to the use of an undefined variable “N”, resulting in a NameError. On the right, the code repeatedly calls the same function due to a logical collapse during generation, eventually exceeding the maximum token limit and leading to a SyntaxError. However, the underlying issue is a latent logical hallucination, rather than the observed syntactic error.
@@ 4 CodeHalu Algorithm
15:                 else
16:                         $\xi\leftarrow\xi\cup\operatorname{State}(\mathsf{GC}_{j}^{\alpha_{i}})$
17:Aggregate and count frequencies of unique $\operatorname{State}(\mathsf{GC}_{j}^{\alpha_{i}})$ in $\xi$
@@
Consider a dataset $\alpha$ contains $k$ samples, where each sample $\alpha_{i}$ consists of a problem description $Q$ and a series of test cases $t_{1},t_{2},\ldots,t_{n}$. Each test case $t_{n}$ includes an input $\mathsf{ip}$ and the corresponding expected output $\mathsf{op}$. Notably, following previous work (Li et al., 2023b; Yan et al., 2023), we integrate resource (time and memory) constraints into the code generation instructions. As shown in Algorithm 1, we use a $\mathsf{\pi}_{j}$ to generate a code solution $\mathsf{GC}_{j}^{\alpha_{i}}$ for each sample $\alpha_{i}$ based on the code generation instruction $\mathsf{GI_{j}}$ and the problem description $\mathsf{Q}$. If $\mathsf{GC}_{j}^{\alpha_{i}}$ exhibits any of the states such as stuttering, infinite loops, or gibberish, we include it in $\xi$.
@@
Finally, we merge the identical states $\operatorname{State}(\mathsf{GC}_{j}^{\alpha_{i}})$ detected by CodeHalu and calculate their occurrence frequencies. According to assumption $\mathsf{ASS}$, $\xi$ can be represented as [($\xi_{1}$, $P_{1}$), …, ($\xi_{o}$, $P_{o}$)], where $\xi_{o}$ denotes the $o^{th}$ type of code hallucination, and $P_{o}$ indicates its frequency. Code hallucinations come from four perspectives: errors, syntax, logic, and execution results. Additionally, CodeHalu is language-agnostic and can adjust to match various programming scenarios depending on the programming language.
5 Code Hallucinations Classification
@@
According to the TIOBE Index [^2] https://www.tiobe.com/tiobe-index/, a metric of programming language popularity, we primarily investigate code hallucinations within the Python. By applying the CodeHalu algorithm on the complex APPS dataset (Hendrycks et al., 2021) and 17 widely-used LLMs, we identify and validate 18 types of hallucination states that violate human expectations during the code generation, including inconsistent code context, ambiguous logic and data flow, conflicting intentions, among others. Using the two-stage heuristic classification method introduced in Remark 8, we categorize code hallucinations into four main types based on the nature and origin of these phenomena: mapping hallucinations, naming hallucinations, resource hallucinations, and logical hallucinations, as illustrated in Figure 1.
@@
Firstly, we extensively reference classification methods for hallucinations in the fields of NLP and multimodal research (Zhang et al., 2023; Ji et al., 2023; Huang et al., 2024; Zhai et al., 2023; Chu et al., 2024), as well as methods for classifying code errors and vulnerabilities in software engineering (Pan et al., 2023; Wang et al., 2024; Huang et al., 2023). We adopt a two-stage heuristic classification strategy. Initially, team members independently review failed code cases and develop preliminary classification frameworks; then, we reach a consensus through collaborative discussions. This widely used approach ensures the adaptability and accuracy of our framework, enabling a systematic understanding of code hallucinations in LLMs.
Secondly, we analyze the cross-task occurrence rates of each model across eight categories of code hallucinations, detailed in Appendix Table 3. The results show that the average cross-task occurrence rate for these categories is only 2.04%, confirming the independence and rationality of our classification. For the Gemma-7B model, which exhibits the most severe hallucinations in Table 2, only 1.07% of task samples show cross-task hallucinations, as illustrated in Figure 3.
Lastly, we conduct an empirical investigation of our classification results and design a questionnaire to evaluate the rationality of our method, detailed in the Appendix Figure 8. The survey receives 23 responses, and after excluding seven respondents with less than three years of experience, we analyzed 16 valid responses. The survey results indicate a rationality rating of 91.08% for our classification method, further supporting its validity.
@@
6 Cause Analysis of Code Hallucinations
Mapping hallucinations stem from the model’s misunderstanding of data types and structures. This phenomenon arises due to several factors: (1) The model generates code based on tokens, lacking insight into higher-level structures such as statements and functions (Yang et al., 2021); (2) When dealing with long-distance data dependencies, especially within complex code blocks, the model fails to continuously track the structure and state of variables, overly relying on local information while neglecting the importance of the overall context (Zhang et al., 2024); (3) The model does not explicitly perform type checking and structure matching during code generation, lacking static checking and error correction mechanisms.
Naming hallucinations reflect the limitations of models in tracking information and utilizing external knowledge. This issue arises from several factors: (1) Token-based feature representation makes it difficult to accurately model long-distance dependencies, leading to model misjudgments regarding variable scope, lifecycle, and visibility (Xu et al., 2020); (2) The code generation process lacks consistency checks for identifiers and does not perform global tracking of variable definitions and usage; (3) Knowledge of external libraries is not effectively and timely integrated into the model’s knowledge system, making it difficult for the model to accurately understand the names, functions, and invocation methods of libraries (Jesse et al., 2023).
Resource hallucinations highlight the model’s lack of deep understanding of code execution mechanisms and physical constraints. These issues arise from several factors: (1) The training data lacks information related to resource consumption and performance optimization, making it difficult for the model to learn about complexity analysis and resource evaluation; (2) As the model generates code based on probabilities, it lacks a module for calculating and estimating the resource consumption of the generated code, making it unable to simulate the real-world operating environment and resource limits; (3) During the model training process, the focus is on the correctness of the code’s functionality, often overlooking its complexity and resource constraints in actual execution environments.
Logic hallucinations reveal the model’s deficiencies in semantic understanding and reasoning about code. This issue arises due to several factors: (1) The model relies on pattern matching and statistical rules to generate code, lacking a fundamental understanding of symbolic systems and rigorous verification of program logic; (2) The training data is often not rigorously verified for accuracy and may contain code with very similar functions. Since models sometimes imitate and memorize previous examples (Yan and Li, 2022), this can result in the model directly replicating similar logic in the code or even learning incorrect logic from the outset; (3) When the model generates code, repetition at the line level has a self-reinforcing effect, causing the model to become increasingly confident in the code it generates, which may lead to a stuttering phenomenon (Xu et al., 2022).
7 The CodeHaluEval Benchmark
We construct the CodeHaluEval benchmark, a unified evaluation method for comparing various types and frequencies of hallucinations in code generation across different LLMs. We develop the CodeHaluEval based on the APPS testing set, following a structured process of Validation-Identification-Construction, as shown in Figure 4.
@@ 7 The CodeHaluEval Benchmark
In the validation phase, we use the CodeHalu algorithm to identify multiple types of hallucinations $\mathsf{HaluTypes}$ $\xi$, represented as [($\xi_{1}$, $P_{1}$), …, ($\xi_{i}$, $P_{i}$)]. In the identification phase, we annotate the ${3k}$ most common hallucinations and their frequencies in each sample $\alpha_{i}$, represented as [($\xi_{1}$, $P_{1}$), …, ($\xi_{3k}$, $P_{3k}$)]. In the construction phase, we sort all samples in descending order based on the frequency $P_{i}$ of each hallucination type $\xi_{i}$. If the hallucination frequency in a sample $\alpha_{i}$ exceeds the threshold $k$, we include this sample in the corresponding hallucination type set in the CodeHaluEval benchmark. When selecting the threshold $k$, we consider both the minimum number of samples required to detect code hallucination effects in the CodeHaluEval benchmark and the inference costs associated with evaluating various LLMs. The efficacy analysis and the selection process for the threshold $k$ are detailed in Appendix A. Through this method, we establish the CodeHaluEval benchmark, with detailed statistics shown in Table 1.
8 Experiments
Models. To comprehensively analyze the different hallucinations of various competitive LLMs in CodeHaluEval, we evaluate 12 general LLMs, including GPT-4 (OpenAI, 2023), GPT-3.5 (OpenAI, 2023), Gemini-Pro-1.0 (Gemini, 2023), Claude-3-haiku (Anthropic, 2024), LLaMA-2 & 3 (Touvron et al., 2023), Vicuna (Chiang et al., 2023), Qwen-turbo (Bai et al., 2023), ChatGLM3-6B (Du et al., 2021), Ernie-3.5(Baidu, 2023), Mistral-7B(Jiang et al., 2023), Gemma (Team et al., 2024). We also evaluate 5 coding LLMs, including Code LLaMA (Roziere et al., 2023), DeepSeek Coder (Guo et al., 2024), CodeGeeX-2 (Zheng et al., 2023), StarCoder-2 (Li et al., 2023a), MagicCoder-7B (Wei et al., 2023), WizardCoder-7B (Luo et al., 2023). The experimental evaluation is conducted using API calls or 8 NVIDIA A6000 GPUs.
@@ 8 Experiments
The experimental results are presented in Table 2 and Figure 5, with case studies of various models across different hallucination categories detailed in Appendix A.
@@
9 Conclusion
We introduce the concept of code hallucination and propose an execution-based verification method to classify code hallucinations. We develop the dynamic detection algorithm, CodeHalu, which categorizes code hallucinations into four main types and eight subtypes, providing a comprehensive understanding of the various challenges faced by LLMs in code generation. Additionally, we establish the CodeHaluEval benchmark and evaluate 17 widely-used LLMs, revealing significant differences in their hallucination patterns during code generation, and providing detailed insights for further improving the code generation capabilities of LLMs. Overall, we lay the theoretical foundation for understanding the hallucination phenomenon of LLMs in code generation, and provide a complete set of tools for detecting and evaluating code hallucinations.
10 Limitations

### v4

@@ Abstract
Introduction
Deep neural networks often generate erroneous information that contradicts the original content, cannot be verified, or conflicts with real-world knowledge. This phenomenon, commonly known as model hallucination, attracts widespread attention in the fields of natural language processing and multimodal learning (Ji et al. 2023; Zhang et al. 2023; Liu et al. 2024), with the community actively exploring methods to mitigate hallucinations (Peng et al. 2023; Elaraby et al. 2023; Liu et al. 2023). However, the issue of model hallucination in the code generation domain remains unexplored.
Conducting a thorough and dedicated study on code hallucinations is crucial for improving the quality of code generated by LLMs. Firstly, the purpose of code is to solve problems, and its value is realized only when the code executes successfully and passes tests (Chen et al. 2021; Austin et al. 2021; Yan et al. 2023). This necessitates that the generated code not only maintain strict logic and precision but also undergoes execution verification to confirm its correctness. Therefore, the practical use and verification of code differ significantly from Natural Language(NL) texts, meaning we cannot directly apply the definitions and methods used for NL hallucinations to code. Secondly, code snippets containing hallucinations may trigger runtime errors, or exhibit functional defects, which hinder the reliable deployment of LLMs in automated software development scenarios. Lastly, by exploring and verifying code hallucinations in a targeted manner, we can effectively uncover their causes and contribute to improving the architecture and training methods of LLMs.
To fill this gap, we define the concept of code hallucination in LLMs, based on the unique purpose and function of the code. Code hallucinations refer to the phenomenon where code generated by LLMs is syntactically correct or even semantically plausible but ultimately cannot execute as expected or fails to meet specified requirements. [^1] We test 16 LLMs using 105,958 code samples. The experimental results demonstrate that only 9 models occasionally exhibit syntactic errors in the generated code, with an exceptionally low average error rate of 0.0020. These findings support our initial hypothesis that the code generated by LLMs is generally syntactically correct and even semantically plausible or appropriate. This phenomenon typically arises from various factors, such as errors or outdated information in the training data, an inadequate grasp of the syntax rules and programming paradigms of the programming languages, and limitations in the logical processing capabilities of the models. In contrast to previous methods that passively explore hallucinations in NLP through a Q&A framework or by prompting LLMs to generate hallucinated answers (Lin, Hilton, and Evans 2021; Cheng et al. 2023), we employ an active strategy to detect hallucinations during the code generation process by LLMs. This approach is crucial as the ultimate goal of the generated code is to execute correctly and fulfill specific tasks.
@@ Introduction
In the field of NLP, hallucination is initially defined as the phenomenon where the text generated by a model is fluent and natural but either lacks substantive meaning or is inconsistent with the provided source content (Ji et al. 2023). Recently, Zhang et al. (2023) standardize the definition of NL hallucinations in LLMs into three categories: input-conflicting hallucinations, where the content generated by LLMs diverges from the user’s input; context-conflicting hallucinations, in which the generated content contradicts previously generated content; and fact-conflicting hallucinations, where the generated content conflicts with established world knowledge. These hallucinations are attributed to various factors, such as poor-quality data samples in the training dataset or the use of sampling algorithms with high uncertainty.
In the multimodal domain, Zhai et al. (2023) classify types of hallucinations in image-to-text scenarios, such as image captioning and visual question answering. They define three main types of hallucinations: object existence hallucinations, object attribute hallucinations, and object relationship hallucinations. In text-to-image scenarios, such as image generation, hallucinations refer to the creation of factually incorrect details by the image generation model in response to the given text input. Huang et al. (2024) introduce VHTest, which evaluates hallucinations across eight dimensions in images, including the existence, shape, color, orientation, OCR, size, position, and counting of visual objects. In text-to-video scenarios, such as video generation, Chu et al. (2024) define three types of hallucinations: prompt consistency hallucinations, static hallucinations, and dynamic hallucinations. Although the issue of hallucinations receives extensive attention in NLP and multimodal domains, it remains unexplored in the code domain. Therefore, we propose CodeHalu to systematically define, identify, classify, and quantify code hallucinations in LLMs.
@@
In recent years, numerous studies focus on evaluating the capability of LLMs to handle various programming tasks. Among these, the HumanEval (Chen et al. 2021), includes 164 Python programming problems, each with an average of 6.7 unit tests. The APPS (Hendrycks et al. 2021) benchmark presents more challenging programming questions, with each problem averaging 293.2 words in length. CodeScope (Yan et al. 2023) covers 43 programming languages and 8 coding tasks to evaluate LLMs in code understanding and generation. MMCode (Li et al. 2024) is designed to evaluate the programming capability of code models in multimodal scenarios. SWE-bench (Jimenez et al. 2023) evaluates the ability of LLMs to modify code repositories and solve problems with a complexity level comparable to what human programmers encounter. Overall, existing code benchmarks focus on evaluating the performance of LLMs on various programming tasks. However, there is still a lack of effective methods to detect and quantify potential hallucinations that may occur in code generation. Therefore, we propose CodeHaluEval to detect and quantify code hallucinations in LLMs.
Code Hallucination
@@
In multiple domains, existing work (Ji et al. 2023; Zhang et al. 2023; Huang et al. 2024; Zhai et al. 2023; Chu et al. 2024) often equates errors with hallucinations, or considers errors as a specific subset of hallucinations. We follow this perspective and regard code errors as a specific subset of code hallucinations. In other words, errors manifest as a form of hallucination, but not all hallucinations can be adequately expressed through errors. Figure 2 illustrates the distinction between code errors and code hallucinations. The code on the left exhibits a typical code error due to the use of an undefined variable “N”, resulting in a NameError. On the right, the code repeatedly calls the same function due to a logical collapse during generation, eventually exceeding the maximum token limit and leading to a SyntaxError. However, the underlying issue is a latent logical hallucination, rather than the observed syntactic error.
@@
15:Aggregate and count frequencies of unique $\operatorname{State}(\mathsf{GC}_{j}^{\alpha_{i}})$ in $\xi$
@@
Consider a dataset $\alpha$ contains $k$ samples, where each sample $\alpha_{i}$ consists of a problem description $Q$ and a series of test cases $t_{1},t_{2},\ldots,t_{n}$. Each test case $t_{n}$ includes an input $\mathsf{ip}$ and the corresponding expected output $\mathsf{op}$. Notably, following previous work (Li et al. 2023b; Yan et al. 2023), we integrate resource (time and memory) constraints into the code generation instructions. As shown in Algorithm 1, we use a $\mathsf{\pi}_{j}$ to generate a code solution $\mathsf{GC}_{j}^{\alpha_{i}}$ for each sample $\alpha_{i}$ based on the code generation instruction $\mathsf{GI_{j}}$ and the problem description $\mathsf{Q}$. If $\mathsf{GC}_{j}^{\alpha_{i}}$ exhibits any of the states such as stuttering, infinite loops, or gibberish, we include it in $\xi$.
@@
Finally, we merge the identical states $\operatorname{State}(\mathsf{GC}_{j}^{\alpha_{i}})$ detected by CodeHalu and calculate their occurrence frequencies. According to assumption $\mathsf{ASS}$, $\xi$ can be represented as [($\xi_{1}$, $P_{1}$), ($\xi_{2}$, $P_{2}$), …, ($\xi_{o}$, $P_{o}$)], where $\xi_{o}$ denotes the $o^{th}$ type of code hallucination, and $P_{o}$ indicates its corresponding frequency. Code hallucinations come from four perspectives: errors, syntax, logic, and execution results. Additionally, CodeHalu is language-agnostic and can dynamically adjust to match various programming scenarios depending on the programming language.
Code Hallucinations Classification
@@
According to the TIOBE Index [^2] https://www.tiobe.com/tiobe-index/, a metric of programming language popularity, we primarily investigate code hallucinations within the Python. By applying the CodeHalu algorithm on the complex APPS dataset (Hendrycks et al. 2021) and 17 widely-used LLMs, we identify and validate 18 types of hallucination states that violate human expectations during the code generation, including inconsistent code context, ambiguous logic and data flow, conflicting intentions, among others. Using the two-stage heuristic classification method introduced in Remark 8, we categorize code hallucinations into four main types based on the nature and origin of these phenomena: mapping hallucinations, naming hallucinations, resource hallucinations, and logical hallucinations, as illustrated in Figure 1.
@@
Firstly, we extensively reference classification methods for hallucinations in the fields of NLP and multimodal research (Zhang et al. 2023; Ji et al. 2023; Huang et al. 2024; Zhai et al. 2023; Chu et al. 2024), as well as methods for classifying code errors and vulnerabilities in software engineering (Pan et al. 2023; Wang et al. 2024; Huang et al. 2023). We adopt a two-stage heuristic classification strategy. Initially, team members independently review failed code cases and develop preliminary classification frameworks; then, we reach a consensus through collaborative discussions. This widely used approach ensures the adaptability and accuracy of our framework, enabling a systematic understanding of code hallucinations in LLMs.
Secondly, we analyze the cross-task occurrence rates of each model across eight categories of code hallucinations. The results show that the average cross-task occurrence rate for these categories is only 2.04%, confirming the independence and rationality of our classification. For the Gemma-7B model, which exhibits the most severe hallucinations in Table 2, only 1.07% of task samples show cross-task hallucinations, as illustrated in Figure 3.
Lastly, we conduct an empirical investigation of our classification results and design a questionnaire to evaluate the rationality of our method . The survey receives 23 responses, and after excluding seven respondents with less than three years of experience, we analyzed 16 valid responses. The survey results indicate a rationality rating of 91.08% for our classification method, further supporting its validity.
@@
Cause Analysis of Code Hallucinations
Mapping hallucinations primarily stem from the model’s misunderstanding of data types and structures. This phenomenon arises due to several factors: (1) The model generates code based on tokens, lacking insight into higher-level structures such as statements and functions (Yang, Liu, and Yin 2021); (2) When dealing with long-distance data dependencies, especially within complex code blocks, the model fails to continuously track the structure and state of variables, overly relying on local information while neglecting the importance of the overall context (Zhang et al. 2024); (3) The model does not explicitly perform type checking and structure matching during code generation, lacking static checking and error correction mechanisms.
Naming hallucinations reflect the limitations of models in tracking information and utilizing external knowledge. This issue arises from several factors: (1) Token-based feature representation makes it difficult to accurately model long-distance dependencies, leading to model misjudgments regarding variable scope, lifecycle, and visibility (Xu et al. 2020); (2) The code generation process lacks consistency checks for identifiers and does not perform global tracking of variable definitions and usage; (3) Knowledge of external libraries is not effectively and timely integrated into the model’s knowledge system, making it difficult for the model to accurately understand the names, functions, and invocation methods of libraries (Jesse et al. 2023).
Resource hallucinations highlight the model’s lack of deep understanding of code execution mechanisms and physical constraints. These issues arise from several factors: (1) The training data lacks information related to resource consumption and performance optimization, making it difficult for the model to learn about complexity analysis and resource estimation; (2) As the model generates code based on probabilities, it lacks a module for calculating and estimating the resource consumption of the generated code, making it unable to simulate the real-world operating environment and resource limits; (3) During the model training process, the focus is usually on the correctness of the code’s functionality, often overlooking its complexity and resource constraints in actual execution environments.
@@
Logic hallucinations reveal the model’s deficiencies in semantic understanding and reasoning about code. This issue arises due to several factors: (1) The model mainly relies on pattern matching and statistical rules to generate code, lacking a fundamental understanding of symbolic systems and rigorous verification of program logic; (2) The training data is often not rigorously verified for accuracy and may contain code with very similar functions. Since models sometimes imitate and memorize previous examples (Yan and Li 2022), this can result in the model directly replicating similar logic in the code or even learning incorrect logic from the outset; (3) When the model generates code, repetition at the line level has a self-reinforcing effect, causing the model to become increasingly confident in the code it generates, which may lead to a stuttering phenomenon (Xu et al. 2022).
The CodeHaluEval Benchmark
We construct the CodeHaluEval benchmark, a unified evaluation method for comparing various types and frequencies of hallucinations in code generation across different LLMs. We develop the CodeHaluEval benchmark based on the APPS testing set, following a structured process of Validation-Identification-Construction, as shown in Figure 4.
@@
In the validation phase, we use the CodeHalu algorithm to identify multiple types of hallucinations $\mathsf{HaluTypes}$ $\xi$, represented as [($\xi_{1}$, $P_{1}$), …, ($\xi_{i}$, $P_{i}$)]. In the identification phase, we annotate the ${k^{2}}$ most common hallucinations and their frequencies in each sample $\alpha_{i}$, represented as [($\xi_{1}$, $P_{1}$), …, ($\xi_{k^{2}}$, $P_{k^{2}}$)]. In the construction phase, we sort all samples in descending order based on the frequency $P_{i}$ of each hallucination type $\xi_{i}$. If the hallucination frequency in a sample $\alpha_{i}$ exceeds the threshold $k$, we include this sample in the corresponding hallucination type set in the CodeHaluEval benchmark. When selecting the threshold $k$, we consider both the minimum number of samples required to detect code hallucination effects in the CodeHaluEval benchmark and the inference costs associated with evaluating various LLMs. Through this method, we establish the CodeHaluEval benchmark, with detailed statistics shown in Table 1.
Experiments
Models. To comprehensively analyze the different hallucinations of various competitive LLMs in CodeHaluEval, we evaluate 12 general LLMs, including GPT-4 (OpenAI 2023), GPT-3.5 (OpenAI 2023), Gemini-Pro-1.0 (Gemini 2023), Claude-3-haiku (Anthropic 2024), LLaMA-2 & 3 (Touvron et al. 2023), Vicuna (Chiang et al. 2023), Qwen-turbo (Bai et al. 2023), ChatGLM3-6B (Du et al. 2021), Ernie-3.5 (Baidu 2023), Mistral-7B (Jiang et al. 2023), Gemma (Team et al. 2024). We also evaluate 5 coding LLMs, including Code LLaMA (Roziere et al. 2023), DeepSeek Coder (Guo et al. 2024), CodeGeeX-2 (Zheng et al. 2023), StarCoder-2 (Li et al. 2023a), MagicCoder-7B (Wei et al. 2023), WizardCoder-7B (Luo et al. 2023). The experimental evaluation is conducted using API calls or 8 NVIDIA A6000 GPUs.
@@
The experimental results are presented in Table 2 & Figure 5.
@@
Conclusion
We introduce the concept of code hallucination and propose an execution-based verification method to classify code hallucinations. We develop the dynamic detection algorithm, CodeHalu, and categorize code hallucinations into four main types, providing a comprehensive understanding of the various challenges faced by LLMs in code generation. Additionally, we establish the CodeHaluEval benchmark and evaluate 17 widely-used LLMs, revealing significant differences in their hallucination patterns during code generation, and providing detailed insights for further improving the code generation capabilities of LLMs. Overall, we lay the theoretical foundation for understanding the hallucination phenomenon of LLMs in code generation, and provide a complete set of tools for detecting and evaluating code hallucinations.
Limitations